OpenStudy (anonymous):

convert 60 degree centigrade into Kelvin

OpenStudy (anonymous):

273+60 = 333

OpenStudy (anonymous):

can u make it more accurate or more exact answer

OpenStudy (anonymous):

i mean in decimals if there would be decimals

OpenStudy (anonymous):

1 kelvin = -272.15 degrees Celsius 272.15 + 60 = 332.15

OpenStudy (anonymous):

\[273.16+60 C = 333.16 K \]

OpenStudy (anonymous):

i had this some where !!!!!!!!!!!!!!!!!!!!!!! is it 273.16 or 272.15 ??????/

OpenStudy (anonymous):

wiki and google says it is 272.15

OpenStudy (anonymous):

and this : http://en.wikipedia.org/wiki/Kelvin

OpenStudy (anonymous):

(273.16 K (0.01 °C; 32.02 °F))

OpenStudy (anonymous):

kk Absolute zero at 0 K is −273.15 °C (−459.67 °F).

OpenStudy (anonymous):

i got it thanks very much

OpenStudy (anonymous):

ypur welcome :))

OpenStudy (anonymous):

r u on facebook ?